The three sides of a triangle measue 20,30 and 40 an the longest side of a similar side is 48, what is the length of the shortest side of this triangle?
Answer by checkley79(3341)

20/X=40/48
40X=20*48
40X=960
X=960/40
X=24 ANS. FOR THE SHORTEST SIDE.